OpenStudy (anonymous):

2 more than the difference of a square of a number and the number gives a second number. The sum of the two numbers is 6. Find the two numbers

OpenStudy (anonymous):

2,4

OpenStudy (anonymous):

the other answer is (-2,8) but can you please show me the translation

OpenStudy (anonymous):

i dont know how to write it correctly

OpenStudy (defiance):

Hello.

OpenStudy (anonymous):

hello i need help with the translation

OpenStudy (defiance):

Hmm.

OpenStudy (defiance):

I really don't know how to explain this either, but the answer is 2 and 4, 2 + 4 = 6.

OpenStudy (anonymous):

the answer(s) are (2,4) , (-2,8)

OpenStudy (defiance):

Well, you could go on for ages with this... -3 + 9, -4 + 10. and such...

OpenStudy (defiance):

ACtually, those wont work

OpenStudy (defiance):

Nvm, :3.

OpenStudy (anonymous):

explain the translation, the teacher is accepting a translation and we are supposed to solve from there

OpenStudy (anonymous):

expecting *

OpenStudy (anonymous):

\[x ^{2}=x\]+2

OpenStudy (anonymous):

X\[x ^{2}= x+2\]

OpenStudy (anonymous):

x+y = 6

OpenStudy (anonymous):

can you write it a bit better its hard to understand

OpenStudy (anonymous):

so u need to make the two conditions happen

OpenStudy (anonymous):

maybe draw it?

OpenStudy (anonymous):

you can make use of this link for plotting equations http://www.wolframalpha.com/

OpenStudy (anonymous):

here is the full answer

OpenStudy (anonymous):

hope this was of any help to u

OpenStudy (anonymous):

this the full answer with the three equations and the plot